OpenStudy (angelwings996):

Simplify the expressin and state any restrictions on a. 4a+3/a+3 * 3a+9/2a-6

OpenStudy (angelwings996):

I know what it is simplified but I am having trouble finding the restrictions

OpenStudy (campbell_st):

ok... so the restiction occur in the denominator of each fraction. you can't have a zero denominator so you need to look at the values that a cannot take. this is done by looking at a + 3 = 0 and 2a - 6 = 0 solving for a will give the values that a cannot take. hope this helps...

OpenStudy (angelwings996):

@campbell_st so I will find the restrictions from the denominator of the two rational expressions above or from my simplified expression ?

jimthompson5910 (jim_thompson5910):

you would look at the original expression and find any restrictions on that (NOT the simplified expression)

OpenStudy (campbell_st):

so looking at the original problem the the restrictions are a = -3 and a = 3 does that make sense...?

jimthompson5910 (jim_thompson5910):

simple example 2x/x = 2 2x/x simplifies to 2 but they are not the same for all values of x unless you explicitly state that x can't be zero

OpenStudy (angelwings996):

@campbell_st Yes that does make sense

OpenStudy (angelwings996):

@jim_thompson5910 Yeah, I understand what you are saying, thanks for the example

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (campbell_st):

because (a + 3) is a common factor still requires it to be included as a restriction. you may learn about the types of restrictions...if you have to graph a similar function a = 3 would be an asymptote and a = -3 would be called a point of discontinuity

OpenStudy (campbell_st):

glad to help

OpenStudy (angelwings996):

Okay so would i say that \[a \neq -3,3\] ?

OpenStudy (campbell_st):

thats correct...

OpenStudy (angelwings996):

Okay thank you!